# Eyes Wide Unshut: Unsupervised Mistake Detection in Egocentric Video by Detecting Unpredictable Gaze

### v2

@@ IV-A1 EPIC-Tent
We base our main experiments on EPIC-Tent [17], a dataset comprising $7$ hours of egocentric video of $29$ subjects wearing two head-mounded cameras (GoPro and SMI eye tracker) while assembling a camping tent. The dataset includes egocentric video, gaze and labels indicating video segments in which users make mistakes. EPIC-Tent contains $151,821$ mistake frames and $384,558$ frames of correct execution, hence with a 28:72 ratio between correct and mistake frames. The unbalanced nature of the dataset arises from the natural rarity of mistakes and makes the problem particularly challenging to address. We randomly split videos in training, validation and test sets roughly following a 60:15:25 ratio.
IV-B Implementation Details
@@
We follow [19] to set the hyperparameters of the GLC module. Differently from [19], we use a stride of $1$, rather than $2$ to train the model, which we found leads to marginal improvements in gaze and mistake prediction. To prevent overfitting, we set the weight decay to 0.07. We set the batch size to $4$ clips of $8$ frames.
IV-C Supervision Levels and Compared Approaches
@@
We compared the proposed approach to methods belonging to two different classes involving different supervision levels: one-class classification methods, and unsupervised methods. Beyond compared methods, we also consider a random baseline in which a random score is assigned to each input clip.
IV-C1 One-Class Classification Methods
@@
are trained only on videos of correct executions, following the standard anomaly detection setup [7, 6]. Note that, while these methods do not access any label, they still require the training set to include only correct executions, hence leading to a one-class setup rather than an unsupervised one. For this class, we compare our method with respect to TrajREC [6] and MoCoDAD [7], two popular approaches for video anomaly detection based on the processing of skeletal data. In the experiment, we replace skeletal data with gaze trajectories. We then consider an instantiation of the proposed approach in which the gaze completion model is trained only on correct executions, hence effectively replicating a one-class paradigm. Our intuition is that the gaze completion module will learn to better model normal behavior, hence leading to more variability in the predictions when a mistake is encountered. Finally, we compare with respect to a baseline which replaces the proposed gaze completion module with a simple gaze prediction component based on GLC [19]. Following the one-class setup, we train the GLC gaze predictor on correct executions and use the investigated scoring functions to predict scores by comparing the ground truth gaze trajectory obtained by the gaze tracker of the wearable device with respect to the predicted gaze.
IV-C2 Unsupervised Methods
@@
assume as input an unlabeled set of clips including both mistakes and correct executions. We compare several variants of the proposed approach to unsupervised mistake detection based on gaze completion, exploring different scoring functions and trajectory conditioning modules, and compare with a gaze prediction baseline which replaces the gaze completion module with GLC and does not condition on partial trajectories.
IV-D Mistake detection on EPIC-Tent
TABLE I: Performance of various methods on EPIC-Tent with respect to the different supervision levels. Per-group best results are reported in bold.
Method Sup. Level Gaze at Test-Time Scoring Fun. AUC F1 Precision Recall
1 Random // 0.51 0.35 0.29 0.42
2 TrajREC [6] One-Class // 0.51 0.45 0.29 0.96
3 MoCoDAD [7] One-Class // 0.50 0.44 0.29 0.97
4 Gaze Prediction One-Class Entropy 0.55 0.41 0.28 0.78
5 Gaze Prediction One-Class Euclidean 0.60 0.41 0.29 0.69
6 Gaze Prediction One-Class DTW 0.61 0.43 0.30 0.74
7 Gaze Prediction One-Class Heatmap 0.66 0.46 0.37 0.62
8 Gaze Completion* One-Class Heatmap 0.66 0.34 0.42 0.29
[…]
11 Gaze Prediction Unsupervised Entropy 0.53 0.42 0.30 0.66
12 Gaze Prediction Unsupervised Euclidean 0.59 0.42 0.32 0.62
13 Gaze Prediction Unsupervised DTW 0.60 0.44 0.33 0.69
14 Gaze Prediction Unsupervised Heatmap 0.61 0.44 0.33 0.70
15 Gaze Completion* Unsupervised Heatmap 0.63 0.45 0.32 0.74
[…]
  • Early Fusion. Mid Fusion. Early and Mid score fusion.
Table I reports the results of the compared mistake detection approaches on the EPIC-Tent dataset, according to the different levels of supervision. For each method, we report the supervision level, whether the approach uses gaze at test time, the scoring function (where applicable) and the evaluation measures. One-class methods originally designed for anomaly detection from skeletal data (lines 2-3) do not improve over the random baseline with respect to the AUC score ($0.50-0.51$ vs $0.51$), and only marginally improve when considering the F1 score ($0.44-0.45$ vs $0.35$). These results suggests that only analyzing gaze trajectories, without relating them to the input video is insufficient for mistake detection. Using a gaze prediction model, paired with the entropy scoring function allows to achieve an AUC of $0.55$ and an F1 of $0.41$. It is worth noting that this approach does not require the availability of gaze at test time. Using different scoring functions which consider gaze as a test-time input allows to improve results in AUC to $0.6$, and $0.61$ when the Euclidean distance and DTW scoring functions are used. The improvement (from $0.55$ of line 4 to $0.61$ of line 6) highlights the usefulness of using gaze as an input signal at test time. Pairing gaze prediction with the heatmap scoring function further boosts the AUC to $0.66$, the F1 score to $0.46$, precision and recall to $0.37$ and $0.62$. These results show the feasibility of comparing predicted and ground truth gaze obtained by the gaze tracker at test time to detect mistakes. Adopting the proposed gaze completion approach allows to further boost results. In particular, this is observed when results of early fusion and mid fusion combined (line 10) [^3] We fuse the two techniques averaging the gaze prediction maps produced by the two models, allowing to achieve an AUC of $0.67$ and and F1 score of $0.47$, the largest score in this supervision level group. While the advantages of gaze completion over gaze prediction are limited in the one-class settings, gaze completion is much more effective than gaze prediction in the unsupervised setting. Indeed, in this case, using gaze completion with early fusion improves results over gaze prediction ($0.63$ of line 15 vs $0.61$ of line 14), and similarly with mid fusion ($0.62$ of line 16 vs $0.61$ of line 14). Our best model (line 17), combining early and mid fusion, achieves an AUC of $0.64$, an F1 score of $0.49$, a precision of $0.34$ and a recall of $0.88$ without seeing any mistake label. It has also a significant improvements over the random baseline (line 1: $0.51$ AUC, $0.35$ F1, $0.29$ precision, $0.42$ recall).
@@ IV-D Mistake detection on EPIC-Tent
TABLE II: Comparison of GLC and GLC with Gaze Completion for Gaze Estimation on Epic-Tent.
Method F1 Recall Precision
GLC 0.37 0.65 0.26
Gaze Completion * 0.40 0.67 0.29
Gaze Completion 0.33 0.63 0.23
  • * Early Fusion. Mid Fusion
IV-E Gaze Estimation results
We have compared the performance of gaze prediction to those of gaze completion across the considered dataset. Table II reports results on the EPIC-Tent dataset. The Gaze Completion with Early Fusion (EF) method, exhibited superior performance, achieving the highest F1 score of 0.40, Recall of 0.67, and Precision of 0.29. This suggests that conditioning on partial trajectories makes the problem of making predictions more structured and hence easier to tackle.
Interestingly, while gaze completion with mid-fusion generally allows to improve over gaze prediction in the mistake detection problem, it does not improve gaze prediction results, suggesting that better mistake prediction is not necessarily achieved by methods able to make better gaze prediction.
@@ IV-E Gaze Estimation results
Figure 4: Qualitative results of gaze completion approaches. Input gaze trajectories are highlighted in orange, predicted trajectories are in green for correct executions and in red for mistakes, ground truth trajectories to be predicted are in blue. Top: examples of correct executions. Our model predicts gaze trajectories (green) coherent with the ground truth (blue). Bottom: gaze is unpredictable when a mistake is made, leading to a large discrepancy between predicted (red) and ground truth (blue) trajectories.
IV-F Qualitative Results
@@
Figure 4 reports qualitative results of the proposed gaze completion approach in the unsupervised setting. For correct executions (Figure 4-top), our model predicts a gaze trajectory (in green) similar to the ground truth one (in blue), taking as input the video and a partial trajectory (in orange). When a mistake is made by the camera wearer (Figure 4-bottom), our algorithm predicts gaze trajectories (in red) significantly different from the ground truth ones (in blue) as gaze becomes unpredictable.
V Conclusion
In this work, we have introduced an approach to detect mistakes in egocentric videos through unsupervised learning, exploiting the gaze signal as input. Grounded in the concept of ”gaze completion”, our approach predicts gaze trajectories based on observed video and partial gaze data. Detection of mistakes is enabled by comparing the predicted trajectory to the ground truth to identify moments in which gaze becomes unpredictable. Experimental validation, conducted on the EPIC-Tent dataset, demonstrates the efficacy of our proposed method, which outperforms existing one-class anomaly detection methods. The results of this study highlight the potential of utilizing gaze information, a readily available data source in many commercial devices, to enhance user experience. By proactively identifying and rectifying errors in task execution, this capability paves the way for more intuitive and interactive wearable technologies that adapt to the user’s needs in real-time.
Acknowledgements
Research at the University of Catania is supported by the project Future Artificial Intelligence Research (FAIR) – PNRR MUR Cod. PE0000013 - CUP: E63C22001940006.

### v3

@@ 4.1 Datasets
EPIC-Tent [13] includes $7$ hours of egocentric video of $29$ subjects wearing a head-mounded GoPro and an SMI eye tracker while assembling a camping tent. The dataset includes egocentric video, gaze and labels indicating video segments in which users make mistakes. EPIC-Tent contains $151,689$ mistake frames and $384,558$ frames of correct execution, hence with a 28:72 ratio between correct and mistake frames, which account to $86,099$, $27,613$, and $37,977$ mistake frames in training, validation, and test sets respectively. The unbalanced nature of the dataset arises from the natural rarity of mistakes and makes the problem particularly challenging to address. We randomly split videos in training, validation and test sets roughly following a 60:15:25 ratio.
IndustReal [26] is designed for studying procedural tasks in industrial-like environments and consists of distinct training, validation, and test sets. The training set comprises $78,902$ frames, with $95.68\%$ frames labeled as correct and $4.32\%$ labeled as mistakes. The validation set includes $38,036$ frames, with $95.18\%$ correct and $4.82\%$ mistaken frames. The test set contains $90,105$ frames, with $92.53\%$ correct and $7.47\%$ mistaken frames.
HoloAssist [36] focuses on a variety of scenarios in which users perform tasks with the assistance of an expert. The training set comprises $11,614,033$ frames, with $94\%$ frames labeled as correct and $6\%$ labeled as mistakes. The test set contains $1,699,562$ frames, with $95\%$ correct and $5\%$ mistaken frames.
4.2 Implementation Details
@@
We base our model on the implementation provided in [16] and set hyperparameters accordingly. We will share the code to provide further details. Differently from [16], we use a stride of $1$, rather than $2$ to train the model, which we found leads to marginal improvements in gaze and mistake prediction. To prevent overfitting, we set the weight decay to $0.07$. We set the batch size to $4$ clips of $8$ frames. Specifically, we process each video with non-overlapping windows of 8 consecutive frames. Each batch contains 4 windows (clips), for a total of 8*4=32 frames per batch.
4.3 Supervision Levels and Compared Approaches
@@
We compare the proposed approach to methods belonging to two different supervision levels: one-class classification and unsupervised. Besides the compared methods, we also consider a random baseline in which a random score is assigned to each input clip.
@@
One-Class Classification Methods are trained only on videos of correct executions, following the standard setup of anomaly detection [7, 32]. Note that, while these methods do not require any label during training, they still require the training set to include only correct executions, hence leading to a one-class setup rather than an unsupervised one. For this class, we compare our method with respect to TrajREC [32] and MoCoDAD [7], two popular approaches for video anomaly detection based on the processing of skeletal data. In the experiments, we replace skeletal data with gaze trajectories. We then consider an instantiation of the proposed approach in which the gaze completion model is trained only on correct executions, hence effectively replicating a one-class paradigm. Our intuition is that the gaze completion module will learn to model normal human behavior, hence leading to discrepant predictions when a mistake is encountered. Finally, we compare with respect to a baseline which replaces the proposed gaze completion module with a simple gaze prediction component based on GLC [16]. Following the one-class setup, we train the GLC gaze predictor on correct executions and use the considered scoring functions to predict scores by comparing the ground truth gaze trajectory obtained by the gaze tracker of the wearable device with respect to the predicted gaze.
@@
Unsupervised Methods area trained on an unlabeled set of clips including both mistakes and correct executions. We evalaute different variants of the proposed approach to unsupervised mistake detection based on gaze completion, exploring different scoring functions and trajectory conditioning modules, and compare with a gaze prediction baseline based on GLC [16].
Table 1: Ablation study results for various scoring functions and fusion strategies on EPIC-Tent in the unsupervised setting. Best results per-block are underlined, while best global results are in bold. CH: channel fusion, CORR: correlation fusion, S: sum fusion.
Scoring Fusion F1 Precision Recall AUC
1 Random // 0.36 0.29 0.42 0.51
2 Entropy // 0.41 0.27 0.62 0.51
3 Euclidean // 0.42 0.29 0.60 0.55
4 DTW // 0.44 0.31 0.68 0.56
5 Heatmap // 0.45 0.32 0.70 0.57
6 Heatmap CH 0.45 0.32 0.74 0.63
7 Heatmap CORR 0.5 0.36 0.82 0.65
8 Heatmap S 0.48 0.35 0.82 0.62
9 Heatmap CH + CORR 0.51 0.36 0.85 0.69
10 Heatmap CH + S 0.49 0.35 0.82 0.65
11 Heatmap CORR + S 0.49 0.35 0.81 0.64
12 Heatmap CH + CORR + S 0.47 0.34 0.8 0.64
4.4 Ablation Study
Table 1 reports an ablation study evaluating the considered scoring functions and gaze-video fusion strategies in the unsupervised mistake detection on EPIC-Tent. In rows 2-5, we compare all scoring functions when all fusion strategies are turned off and the model is not conditioned on previous trajectories. The entropy scoring function achieved an AUC of 0.51 and an F1 score of 0.41 (row 1), only marginally above the random baseline, leading to the conclusion that high entropy in the prediction only partially correlates with the presence of a mistake. Notably, this approach does not require gaze data at test time. Employing alternative scoring functions that consider gaze as a test-time input improved the results, yielding an AUC of 0.55 and 0.56 when using Euclidean distance and DTW scoring functions, respectively (rows 2-3). The heatmap-based scoring function produced the best results in this comparison, with an AUC of 0.57 and an F1 score of 0.45, significantly above random level (compare with the F1 score of $0.36$ in row 1). The advantage of the random baseline is likely due to the better exploitation of the probability values computed by the gaze prediction model, as compared to other scoring functions. Hence, we adopted the heatmap-based scoring as our primary method in following comparisons. Rows 6-8 compare approaches using one of the three fusion strategies. While all fusion strategies improve results (compare row 6-8 with 5), the proposed correlation strategy (CORR) systematically outperforms channel and sum fusion, obtaining an AUC score of $0.65$ and an F1 score of $0.5$ and doubling the recall of the random baseline ($0.82$ vs $0.42$) with a better precision ($0.36$ vs $0.29$). Rows 9-12 finally explore combinations of fusion strategies. The best results are obtained by combining our proposed Correlation fusion with Channel fusion, achieving an AUC of 0.69 and an F1 score of 0.51 (+0.12 and +0.06 compared to the standard heatmap method - row 5). This configuration is the one referred to as “ours” in future comparisons, unless stated otherwise. Interestingly, adding the Sum fusion strategy leads to worse results. We speculate that this is due to the challenges associated with sum fusion, which requires an expressive mechanism to map trajectory information to the mid-level representations of the transformer encoder. Moreover, we can see the correlation strategy as both an early fusion method (at the token embedding layer) and as a mid fusion method (at the correlation level), which likely allows correlation fusion to capture similar information as sum fusion.
@@ 4.4 Ablation Study
Table 2: Comparison of GLC and the proposed Gaze Completion approach for Gaze Estimation on EPIC-Tent.
Method Fusion F1 Recall Precision
GLC // 0.37 0.65 0.26
Gaze Completion CH 0.40 0.67 0.29
Gaze Completion CH + CORR 0.43 0.70 0.31
Table 2 compares the performance of GLC [16] used for standard gaze prediction with respect to different versions of our gaze completion approach on the EPIC-Tent dataset. Just using channel fusion brings a performance boost, achieving an F1-score of 0.40, recall of 0.67, and precision of 0.29, while combining channel and correlation fusion brings best results with an F1-score of 0.43, recall of 0.70, and precision of 0.31. These results suggest that conditioning on partial trajectories makes the problem of gaze prediction less uncertain and the proposed approach can leverage the informative content provided by the input trajectory surpassing the performance of a standard gaze prediction method. Moreover, performance in Table 2 correlates with performance in Table 1, suggesting that good gaze prediction is essential to mistake detection performance.
Correlation between self-reported uncertainty and mistakes: Our main assumption is that gaze patterns can signal the presence of a mistake due to the user feeling disoriented and exhibiting abnormal attention patterns. We conducted a logistic regression analysis to explore the relationship between self-reported uncertainty levels (normalized in a $1-5$ range) included in the EPIC-Tent dataset [13] and the occurrence of labeled mistakes. Our analysis revealed that an increase in self-reported uncertainty by 1 unit is associated with a doubling the odds of the occurrence of a mistake  [^4] We obtain a coefficient of 0.68 with a $p<0.001$.. Figure 5 illustrates the distributions of self-reported uncertainty values with respect to the occurrence of mistakes.
Table 3: Mistake detection result on EPIC-Tent. Best results are in bold.
Method Sup. Level F1 Precision Recall AUC
Random // 0.36 0.29 0.42 0.51
TrajREC [32] One-Class 0.45 0.29 0.96 0.51
MoCoDAD [7] One-Class 0.44 0.29 0.97 0.50
GLC [16] One-Class 0.46 0.37 0.62 0.66
Ours One-Class 0.52 0.37 0.85 0.69
GLC [16] Unsupervised 0.44 0.33 0.70 0.61
Ours Unsupervised 0.51 0.36 0.85 0.69
4.5 Comparison with Competitors
Table 3 compares the proposed approach with competitors on the EPIC-Tent, according to the two considered levels of supervision. One-class methods such as TrajREC and MoCoDAD, originally designed for anomaly detection from skeletal data, show minor improvements over the random baseline in terms of AUC score ($0.50-0.51$ vs $0.51$) and only small improvement in F1 score ($0.44-0.45$ vs $0.35$). High recall values, paired with low precision, suggest that such methods tend to classify most clips as mistakes. These results indicate that solely analyzing gaze trajectories without considering the input video is insufficient for effective mistake detection. The state-of-the-art gaze prediction approach, GLC [16], shows improved results compared to the one-class methods, highlighting the advantages of incorporating gaze analysis for mistake detection. The proposed method brings best results yielding an AUC of 0.69, F1-score of 0.52, precision of 0.37, and recall of 0.85, improvements of $+3/6\%$ with respect to the best approach GLC and $+10/20\%$ with respect to the random baseline. While the benefits of gaze completion over gaze prediction are small but systematic in one-class settings, gaze completion proves significantly more effective in unsupervised scenarios. Indeed, in this case, our method brings significant improvements over both the random baseline (e.g., $0.69$ vs $0.51$ AUC score and $0.51$ vs $0.36$ F1 score) and the GLC competitor with $+8\%$ in AUC score, $+7\%$ in F1 score, $+3\%$ in precision and $+15\%$ in recall.
Table 4: Mistake detection result on HoloAssist.
Method Sup. Level F1 Precision Recall AUC
Random // 0.04 0.02 0.4 0.5
TrajREC [32] One-Class 0.09 0.04 0.96 0.5
MoCoDAD [7] One-Class 0.11 0.06 0.94 0.51
GLC [16] One-Class 0.19 0.11 0.56 0.60
Ours One-Class 0.22 0.14 0.59 0.61
GLC [16] Unsupervised 0.10 0.06 0.34 0.54
Ours Unsupervised 0.18 0.12 0.40 0.59
Table 5: Mistake detection result on IndustReal.
Method Sup. Level F1 Recall Precision AUC
Random // 0.12 0.06 0.62 0.51
TrajREC [32] One-Class 0.18 0.10 0.90 0.56
MoCoDAD [7] One-Class 0.19 0.11 0.91 0.57
GLC [16] One-Class 0.21 0.15 0.33 0.60
Ours One-Class 0.24 0.18 0.35 0.63
GLC [16] Unsupervised 0.21 0.15 0.33 0.58
Ours Unsupervised 0.21 0.16 0.33 0.62
Figure 4: Uncertainty boxplot
EPIC-Tent
IndustReal
HoloAssist
Figure 5: Failure cases by action classes
Table 4 compares the proposed method with competitors on HoloAssist. The HoloAssist dataset presents a more varied and expansive context compared to EPIC-Tent, including different sorts of activties, which makes it significantly more challenging to perform mistake detection. This is evidenced by the fact that the Random baseline achieves an F1-score of only $0.04$. One-class baselines like TrajREC and MoCoDAD bring improvement over the baseline but suffer from a tendency to overly classify clips as mistakes (Recall $0.96$ and $0.94$). GLC brings significant improvements in AUC and F1 scores, with slighlty more balanced precision and recall values. Best results are obtained by our approach, which achieves an AUC of 0.61 and an F1-score of 0.22 in one-class settings. Performances are slightly lower in the unsupervised case with an AUC score of $0.59$ and an F1 score of $0.18$, which still exhibit significant improvements over GLC, with gains of $+5\%$ and $+8\%$ respectively.
@@ 4.5 Comparison with Competitors
Figure 6: Qualitative results of the proposed approach. Input gaze trajectories are highlighted in orange, predicted trajectories are in green for correct executions and in red for mistakes, ground truth trajectories are in blue.
Table 5 reports results on the IndustReal dataset. Results confirm previous observations with TrajREC and MoCoDAD bringing minor improvements over the random baseline, and our method outperforming all competitors and improving over GLC by systematic margings both the one-class case ($+3\%$ in AUC and $+4\%$ in F1) as well as in the unsupervised one ($+4\%$ in AUC score and comparable F1 performance).
4.6 Qualitative Results and Failure Cases
@@
Figure 6 reports qualitative results of the proposed gaze completion approach in the unsupervised setting. For correct executions (Figure 6-top), our model predicts a gaze trajectory (in green) similar to the ground truth one (in blue), taking as input the video and a partial trajectory (in orange). When a mistake is made by the camera wearer (Figure 6-bottom), our algorithm predicts gaze trajectories (in red) significantly different from the ground truth ones (in blue).
Figure 5 reports failure cases by action classes in EPIC-Tent (a), IndustReal (b) and HoloAssist (c). Most failures are with actions requiring fine eye-hand coordination (insert support, align) or actions not involving interactions (instruction, inspect), few failures are observed in simple actions (pick up/open/place, take, place).
5 Conclusion
In this work, we introduced an approach for detecting mistakes in egocentric videos in an unsupervised way, leveraging gaze signals as input. We proposed a novel gaze completion task, where we predict gaze trajectories based on observed video and partial gaze data and an approach to tackle this task. Detection of mistakes is performed by comparing predicted trajectories against ground truth, identifying instances where gaze behavior becomes unpredictable as potential mistakes. Experimental validation on the EPIC-Tent, HoloAssist, and IndustReal datasets demonstrates the efficacy of our method, surpassing traditional one-class anomaly detection techniques and other unsupervised mistake detection methods. Our method is ranked first on the HoloAssist Mistake Detection challenge.
Limitations. Without strong assumptions on the semantics of the task, the proposed method is able to detect certain classes of mistakes, mainly related to fine hand-object coordination, while it may miss other classes of mistakes such as those related to key-step ordering for which procedure-aware methods may be beneficial.